OpenStudy (anonymous):

What are the solutions of 5x2 = -15x?

OpenStudy (anonymous):

\[\LARGE 5x^2=-15x\] \[\LARGE 5x^2+15x=0\] .. note that here you have , the quadratic formula: use its formula: \[\LARGE x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] you have: \[\LARGE 5x^2+15x+0=0\] now substitute a=5 b=15 c=0 and there you go :)

OpenStudy (anonymous):

thanks